OpenStudy (18jonea):

Determine the number and type of complex solutions and possible real solutions for each of the following equations. 2) 4x^3 – 12x + 9 = 0 @mathmale @michele_laino

OpenStudy (18jonea):

Would I start by doing b^2 -4ac?

OpenStudy (michele_laino):

no, since it is a third degree equation

OpenStudy (18jonea):

ok so what would i do?

OpenStudy (michele_laino):

I'm thinking...

OpenStudy (fibonaccichick666):

can you tell us how many solutions you will have? There is a theorem you will have learned that will tell you this.

OpenStudy (albert0898):

Well, you have a few options. 1) You can plug it into your calculator and find how many zeros there are. 2) You can use the p/q method. There should be a few more

OpenStudy (18jonea):

Wouldnt there be 3 solutions since it is to the third degree?

OpenStudy (michele_laino):

yes! such equations has three solutions

OpenStudy (michele_laino):

equation*

OpenStudy (michele_laino):

it is suffice to apply the fundamental theorem of algebra

OpenStudy (18jonea):

ok so what do i do to determine the number and type of complex solutions and possible real solutions for each of the following equations.

OpenStudy (18jonea):

I have to be able to show work

OpenStudy (fibonaccichick666):

So, the FTA tells us how many roots there are. First thing we want to do for work is try and factor this into something nice.

OpenStudy (fibonaccichick666):

so let's try and reduce this to some sort of multiplication we already know how to find roots of.

OpenStudy (18jonea):

Ok how woild we do that

OpenStudy (fibonaccichick666):

play with it algebraically. should be able to get (x+__)(quadratic)=0

OpenStudy (18jonea):

I honestly have no idea

OpenStudy (fibonaccichick666):

but check your op values first because the numbers for this are really really nasty

OpenStudy (fibonaccichick666):

for your reference of how nasty they are http://www.wolframalpha.com/input/?i=+4%28x%5E3+ –+3x%29+%2B+9+%3D+0

OpenStudy (fibonaccichick666):

oh wait, we don't actually have to find the roots just say what they are?

OpenStudy (18jonea):

Determine the number and type of complex solutions and possible real solutions for each of the following equations.

OpenStudy (18jonea):

so it doesnt say we have to find them

OpenStudy (michele_laino):

I think it is better to go into the complex plane \(z\), namely, I rewrite your equation, like below: \[\Large 4{z^3} - 12z + 9 = 0,\quad z = x + iy\]

OpenStudy (fibonaccichick666):

so, by FTA how many solutions do we have to have?

OpenStudy (fibonaccichick666):

then, when you graph it in the real plane, how many zeroes are there?

OpenStudy (fibonaccichick666):

if only 1, we know that there must be a complex root and those only come in pairs.

OpenStudy (fibonaccichick666):

I think that'd be my approach.

OpenStudy (18jonea):

is the 0 where x= 0

OpenStudy (zarkon):

descarte's rule of signs tells us there is exactly one negative real zero

OpenStudy (michele_laino):

if I multiply both sides by the cnjugate \({\bar z}\), I get: \[\Large 4{z^2}{\left| z \right|^2} - 12{\left| z \right|^2} + 9\bar z = 0\]

OpenStudy (michele_laino):

conjugate*

OpenStudy (zarkon):

then using calculus one can show that the negative root is the only real one...thus the rest are complex

OpenStudy (18jonea):

Ok but somehow i have to show my work with an equation

OpenStudy (18jonea):

So how ould i do that

OpenStudy (michele_laino):

next I replace these quantities: \[\Large \begin{gathered} {z^2} = {x^2} - {y^2} + 2ixy \hfill \\ {\left| z \right|^2} = {x^2} + {y^2} \hfill \\ \bar z = x - iy \hfill \\ \end{gathered} \]

OpenStudy (18jonea):

We have never used the z one either

OpenStudy (18jonea):

@mathmale

OpenStudy (18jonea):

?

OpenStudy (michele_laino):

sorry, my procedure, doesn't work

OpenStudy (18jonea):

do you know any other way

OpenStudy (michele_laino):

I'm searching for it, please wait...

OpenStudy (18jonea):

@mathmale

OpenStudy (michele_laino):

we can compute the real solution easily, if we solve graphically, this algebraic system: \[\left\{ \begin{gathered} y = 4{x^3} \hfill \\ y = 12x - 9 \hfill \\ \end{gathered} \right.\] here is the solution:

OpenStudy (18jonea):

I dont think that is the real solutionis it?

OpenStudy (michele_laino):

as we can see, we have only one real solution

OpenStudy (18jonea):

There has to be one way where we can find b0oth real and complex i just dont know what that equation would be

OpenStudy (18jonea):

@mathmale do you have any idea

OpenStudy (18jonea):

actually the real solution is right but how do i show how to find the complex>?

OpenStudy (18jonea):

?

OpenStudy (18jonea):

@triciaal

OpenStudy (triciaal):

sorry

OpenStudy (triciaal):